A piechart represents the results of an election on an alien planet. The candidate, Hum, won the election, receiving five times as many votes as Yab. Pip received 30% of all votes. What percentage of votes did Hum receive?
Answer by checkley71(8403)

PIP GOT 30% OF THE VOTES.
THEREFORE THAT LEAVES 70% OF THE TOTAL VOTES.
5X+X=.70
6X=.70
X=.70/6
X=.11 2/3 OR 11 2/3% WAS THE AMOUNT OF YAB'S VOTES.
70-11 2/3=58 1/3% WAS THE AMOUNT OF hUM'S VOTES.
PROOF:
30+11 2/3+58 1/3=100
100=100